# A Two-Stage Stochastic Model for Road-Rail Intermodal Freight Transportation Under Demand and Capacity Uncertainty

### v1

@@ 1 Introduction
Intermodal freight transportation involves the movement of freight using at least two transportation modes without changing the load unit throughout the process. This approach leverages the cost-effectiveness and high capacity of modes like rail or sea for long-haul segments while relying on trucks for shorter distances. It typically requires consolidating multiple shipments into containers or single units at facilities such as warehouses or manufacturing points. These consolidated units are then transported to intermodal hubs for shipping and eventually transferred to final intermodal nodes for deconsolidation and final delivery. The scale and importance of intermodal freight are evident in the fact that, in 2021, the United States (US) alone transported an average of 53.6 million tons of freight daily, valued at over \$54 billion (Bureau of Transportation Statistics,, 2021), emphasizing it’s significant role in the national economy.
Despite its advantages, intermodal freight transport faces several challenges due to the involvement of multiple stakeholders and decision-makers, including shippers, carriers, and policymakers, each with their own priorities and objectives. In the context of our work, a significant challenge is the modal shift from road to rail transport (El Yaagoubi et al.,, 2022). Integrating these modes efficiently requires overcoming additional handling and time requirements compared to direct road transport (Görçün et al.,, 2024). For longer distances, rail’s high fixed costs are offset by increased efficiency, while road transport becomes less cost-effective as distance increases (Zgonc et al.,, 2019).
@@
Another critical challenge is managing the uncertainty in market demand and spot capacities at intermodal hubs. Demand fluctuations can arise from various factors, including booking changes, cancellations, unforeseen events, and seasonal variations (Wang and Meng,, 2021). Compounding this uncertainty is the stochastic nature of spot capacities, which depend not only on competing cargo from other companies but also on the variability of loading and unloading activities at each station (Sun and Schonfeld,, 2016). These dynamics create a complex environment where maintaining operational efficiency requires careful planning and adaptability.
@@
A key aspect influencing spot capacity allocation is the role of long-term contracts. These agreements are typically made between transport providers and large companies or freight forwarders, ensuring a guaranteed capacity over a specified period. Such contracts offer stable and predictable volumes for the transport providers while securing regular service for the customers. However, any unused capacity from these agreements is often redirected to the spot market. This cascading allocation process introduces additional uncertainty for spot customers, as the availability of capacity depends on how much of the allocated slots long-term contract holders utilize (Wang and Meng,, 2021). For instance, if contract holders underutilize their guaranteed slots, the surplus is released for spot demand, but this availability may still be constrained by the need to accommodate delayed or overbooked cargo from other services or other long-term contractors. This interplay between long-term contracts and spot demand underscores the need for flexible and adaptive strategies to ensure optimal utilization of available capacity while addressing the inherent uncertainty in both supply and demand (Lu et al.,, 2020; Delbart et al.,, 2021).
In this study, we address the problem faced by manufacturing and logistics companies in managing intermodal transport operations with predefined and fixed train schedules, specifically focusing on container transport. Our primary emphasis is on the critical modal shift from road to rail, highlighting the need for precise timing and coordination to meet fixed train schedules efficiently and reliably. Containers are transported from multiple warehouses or manufacturing plants to intermodal hubs or train stations using trucks. However, we do not consider movements beyond the final train stations, as these often serve as centralized distribution hubs with independent local logistics networks (Rodrigue,, 2020). By narrowing the scope to the intermodal portion of the supply chain, we aim to optimize this resource-intensive segment, which offers significant opportunities for efficiency and environmental benefits.
This problem is further complicated by the stochastic nature of demand and uncertainties in spot capacities at each station, as highlighted earlier. Companies must determine the optimal number of containers to prepare at supply nodes, select facilities for container transport, and identify suitable intermodal hubs or train stations for delivery, all while accounting for variability in spot capacity availability (see Fig. 1). The complexity is heightened by multiple trains with fixed schedules and the cascading effects of unused long-term contract capacities on spot market availability. Our approach incorporates sustainability objectives by minimizing carbon emissions during transportation, enhancing both operational efficiency and environmental performance.
We propose a two-stage stochastic optimization model that integrates carbon reduction strategies. The objective of the model is to minimize a mean-risk function that accounts for various components of the total random cost. These include the operational cost of preparing containers at the origins, transportation costs from origins to intermodal hubs or train stations, penalties for unmet demand, and penalties for exceeding carbon emissions limits. To address the uncertainty inherent in demand and available capacity at each station, our model incorporates Conditional Value-at-Risk (CVaR) as part of the objective function. CVaR is a widely-used risk measure that focuses on minimizing the potential for extreme losses. It accounts for the “tail risk" by capturing the expected losses exceeding a specified quantile in the upper tail of the loss distribution. In our case, CVaR helps balance cost efficiency with risk management by emphasizing the potential downside risks associated with demand fluctuations and capacity uncertainties.
In the first stage, the model determines the optimal supply levels at each origin and the initial transportation of containers to train stations before the uncertainty is realized. As for the second stage, after the uncertainty is realized, the model re-optimizes the allocation of containers to meet the demand as closely as possible utilizing the available capacity while minimizing penalties for unmet demand and excess emissions. The model ensures the transportation network adapts to actual conditions, balancing the trade-offs between demand and capacity requirements, costs, and operational efficiency under stochasticity.
To showcase the practical application of our model, we incorporate a case study derived from a segment of the extensive rail network in the US. This case study examines the integration of road and rail transportation, focusing on five critical freight train routes that link a total of 26 intermodal hubs with 68 FedEx warehouses distributed across multiple states, ensuring a minimum of two warehouses per state. Utilizing real-world data on train schedules, distances, and facility locations from the Freight Analysis Framework (FAF) and the National Transportation Atlas Database (NTAD), the study offers a comprehensive view of the operational dynamics within the network.
@@ 2 Literature review
Intermodal freight transport has been widely studied for its potential to improve logistics efficiency, sustainability, and resilience. Recent advancements in optimization techniques, data-driven decision-making, and sustainable transport policies have addressed key challenges such as managing uncertainties, coordinating operations across modes, and minimizing emissions. This section highlights significant recent contributions in these areas, examines emerging approaches, and identifies research gaps, providing context for the proposed model’s contributions.
@@ 2.1 Managing Uncertainties in Intermodal Transport
Managing uncertainties in intermodal transport has been a focal point for researchers aiming to enhance the resilience and efficiency of logistics systems. Guo et al., (2020) tackled dynamic and stochastic shipment allocation challenges by proposing a multistage stochastic programming model. Their anticipatory optimization approach effectively integrated rolling horizon techniques to manage future demand fluctuations, outperforming traditional methods in reducing costs and improving allocation efficiency. Similarly, Sun, (2020) employed a fuzzy multi-objective routing model to address demand uncertainties in hazardous materials transportation. By incorporating triangular fuzzy numbers and soft time windows, their model balanced risk, economic considerations, and service reliability, providing a comprehensive solution for uncertainty in freight transport.
Facility location optimization and disruption management have also been critical to addressing transport uncertainties. Badyal et al., (2023) introduced a two-stage stochastic model to optimize terminal locations and freight distribution in networks prone to disruptions. Their findings emphasized the need to integrate robust planning to enhance cost and resilience. Extending this focus, Bagheri et al., (2024) compared deterministic, stochastic, and robust approaches for optimizing dry port locations. Their study highlighted the superior adaptability of robust models to dynamic conditions, showcasing their potential to reduce environmental impacts while maintaining operational efficiency.
Routing and network optimization remain pivotal in mitigating operational uncertainties. Lu et al., (2020) introduced a fuzzy programming model to manage intercontinental road-rail transport uncertainties, achieving improved efficiency through sensitivity analyses and defuzzification techniques. Yang et al., (2024) expanded on routing challenges by developing a bi-level genetic algorithm to optimize multimodal container transport paths, addressing time-sensitive and general cargo delivery constraints. This method proved effective in minimizing costs and ensuring timely deliveries, offering a scalable solution for complex logistics systems. Wang et al., (2024) complemented these efforts with a constrained optimization model designed to address marine transport delays, employing adaptive differential evolution to significantly reduce costs and delays in multimodal networks.
Integrating demand and capacity uncertainties within broader network designs has yielded impactful results. Li et al., (2023) proposed a multi-commodity flow model using Benders decomposition to optimize shared freight and passenger railway capacity, achieving a balance between operational costs and user satisfaction. Wang et al., (2022) introduced a tri-level programming model with a Kriging-surrogate-based algorithm to optimize large-scale multimodal networks. Their innovative approach demonstrated significant computational efficiency and scalability, solving real-world networks with over 120,000 origin-destination pairs.
Lastly, decision-making frameworks have played a key role in addressing systemic uncertainties and stakeholder priorities. Gandhi et al., (2024) employed a hybrid multi-criteria decision-making framework to evaluate modal shift risks and enablers, emphasizing the need for long-term partnerships and policy interventions. Raad and Rajendran, (2024) further contributed by developing a hybrid fuzzy stochastic model to design closed-loop dry port networks, enhancing resilience through scenario-based robust optimization. These frameworks underline the importance of integrating strategic decision-making tools to build sustainable and adaptive intermodal logistics networks.
@@ 2.2 Operational Synchronization in Road-Rail Intermodal Transport
Operational synchronization between road and rail is crucial for reducing delays and improving intermodal efficiency. Newman and Yano, (2000) pioneered train schedule optimization using decomposition techniques to streamline scheduling and meet delivery requirements. Kravets et al., (2021) introduced a multicriteria model to minimize transit times, demonstrating efficiency gains by consolidating cargo at rail stations. Liu et al., (2024) addressed intermodal port challenges with a flexible yard crane scheduling model, optimizing task sequencing and reducing truck-induced delays, particularly in high railway container ratio environments.
Scheduling and routing optimization have further contributed to improving the performance of intermodal networks. Chen et al., (2016) examined uncertainties in vehicle scheduling and dispatching at intermodal terminals, demonstrating significant improvements in logistics coordination and reduction in delay. Meanwhile, Chen et al., (2022) developed a bi-level optimization model integrating route planning and timetable adjustments, achieving high capacity utilization and reduced costs in China-Europe rail networks. Additionally, Tong et al., (2024) focused on emergency response scenarios, presenting a hybrid flow shop scheduling model and a memetic algorithm to minimize congestion and transshipment delays in road-rail multimodal systems. These studies underscore the importance of tailored scheduling solutions to address specific operational challenges in intermodal logistics.
Strategic planning and infrastructure improvements are essential for synchronized intermodal operations. Delgado et al., (2021) developed an optimization framework for decentralized terminal management, reducing costs and improving terminal planning in Portugal’s rail-road network. Li et al., (2022) tackled bottlenecks in China-Europe freight transport with a network equilibrium model that identified critical transfer points and guided infrastructure expansion. In emerging economies, Kumar and Anbanandam, (2020) highlighted poor infrastructure and limited transshipment technologies as major barriers, recommending targeted investments and policy interventions to enhance intermodal efficiency.
The resilience of intermodal systems to disruptions is another critical aspect of synchronization. Pasquale et al., (2022) proposed a two-stage modeling approach to analyze the ripple effects of disruptions, such as railway arc failures, on passenger and freight flows. Their simulations revealed key dependencies within intermodal networks, informing resilience strategies. Ke and Verma, (2021) extended this focus by developing a framework to manage disruption risks in rail-truck intermodal networks, employing predictive analytics to identify critical terminals and optimize recovery strategies. Additionally, Guo et al., (2024) developed International Multimodal Transport Connectivity (IMTC) index, which assessed transport performance and highlighted the resilience of the China Railway Express (CR Express) during COVID-19 disruptions. These findings underscore the importance of resilient design in ensuring seamless intermodal operations.
Finally, addressing systemic barriers and operational inefficiencies is essential for improving synchronization. Ke, (2022) explored robust optimization strategies for hazardous material transportation, focusing on rerouting, yard repairs, and third-party carrier utilization under yard disruptions. By enhancing network reliability and mitigating disruptions, their work highlighted the value of resilience-focused strategies. Similarly, Ahmady and Eftekhari Yeghaneh, (2022) developed a multi-objective optimization model that balances transport costs and system efficiency while addressing the criticality of infrastructure components, contributing to the broader goal of synchronized operations.
@@ 2.3 Decarbonization in Intermodal Freight Transport
The pursuit of decarbonization in intermodal freight transport has spurred innovative strategies leveraging optimization models and policy interventions. Cheng and Tan, (2019) developed a multimodal path optimization model that minimized carbon emissions and costs by incorporating carbon taxation and trading within time-constrained freight systems. Their genetic algorithm-based approach demonstrated the feasibility of achieving low-carbon transport strategies while maintaining operational timeliness. Similarly, Sun et al., (2023) introduced a fuzzy nonlinear programming model to optimize carbon-efficient road-rail intermodal routing, balancing environmental and operational goals through time-dependent and uncertain transport conditions.
Technological advancements and fuel alternatives have emerged as critical drivers for reducing emissions in intermodal transport. Hernandez et al., (2024) presented the NUFRIEND framework to evaluate alternative fuel technologies for freight rail decarbonization, focusing on biodiesel, e-fuels, hydrogen, and battery-electric systems. Their life-cycle and techno-economic analyses highlighted battery-electric locomotives as a viable solution, achieving a 46% reduction in emissions under specific cost conditions. Matsuyama et al., (2024) echoed the importance of technological innovation in Japan’s hinterland transport system, demonstrating that advanced technologies could drive emission reductions of up to 87%, surpassing the impacts of modal shifts.
Policy-driven modal shifts also play a pivotal role in reducing the carbon footprint of intermodal systems. Halim, (2023) evaluated the impact of rail freight enhancement policies in Indonesia, revealing that combined fiscal measures and infrastructure investments could increase rail freight’s modal share to 20% by 2030, leading to a 21.6% reduction in CO2 emissions. Similarly, Nassar et al., (2023) used a System Dynamics model to analyze the effectiveness of modal shift policies in Brazil, emphasizing the need for infrastructure investments alongside fiscal and regulatory measures to accelerate decarbonization efforts. These studies underline the necessity of integrating multimodal strategies with policy frameworks to achieve sustainable freight transport.
Efforts to align intermodal strategies with emission reduction goals have also focused on collaborative and game-theoretic approaches. Zhang et al., (2022) proposed a collaborative planning framework that optimized capacity utilization and reduced emissions by incorporating eco-label preferences in intermodal freight transport. In parallel, Tamannaei et al., (2021) used a Stackelberg game model to evaluate government interventions in road-to-rail modal shifts, finding that well-designed fuel taxes and subsidies could enhance sustainability metrics without significant economic drawbacks. These findings emphasize the value of coordinated efforts between stakeholders and policymakers to foster sustainable intermodal operations.
Incorporating consolidation strategies into intermodal networks has further enhanced decarbonization efforts. Goodarzi et al., (2024) proposed a two-stage stochastic model integrating environmental and social metrics to address disruption scenarios effectively. Their results demonstrated substantial cost savings and resilience improvements, highlighting the potential of consolidation investments to yield significant environmental and economic benefits. Similarly, López-Acevedo et al., (2024) showcased the feasibility of short-distance intermodal solutions, where strategically located terminals optimized rail usage, reducing emissions and transport costs in the Madrid region.
Together, these studies illustrate the multifaceted approaches required to decarbonize intermodal freight systems, emphasizing the importance of integrating technological advancements, policy-driven modal shifts, and strategic network optimizations. The collective insights from these works provide a robust foundation for sustainable freight transport planning that balances environmental and operational goals.
@@ 2.4 Our Contribution
Building on the advances in the literature, our work addresses critical gaps in managing uncertainties, operational synchronization, and decarbonization within intermodal freight transport systems. We propose a two-stage stochastic optimization model tailored for road-rail intermodal freight transport that addresses critical gaps in literature. Unlike previous works that focus on isolated aspects such as shipment allocation under demand uncertainties (Guo et al.,, 2020) or carbon-efficient routing (Sun et al.,, 2023), our model integrates these components within a unified framework. By leveraging CVaR to account for extreme cost scenarios, our approach provides a robust mechanism to manage uncertainties while balancing cost efficiency and environmental sustainability.
A key aspect of our work is its incorporation of real-time adaptive responses to stochastic elements such as train capacities and demand fluctuations through the utilization of spot capacities at intermodal hubs. This dynamic approach ensures that decisions regarding container preparation and allocation aligns with operational constraints while optimizing resource utilization to meet demand efficiently. In contrast to static models Badyal et al., (2023), which focus primarily on long-term facility location planning, our model actively integrates recourse actions, enabling responsive adjustments to narrow time windows and fluctuating capacities at intermodal hubs.
Another distinctive feature of our model is its dual emphasis on minimizing costs and reducing environmental impact. While prior studies such as Hernandez et al., (2024) have evaluated the potential of alternative fuels, our model embeds emissions constraints directly into the optimization framework. By penalizing excess emissions and incorporating these penalties into the objective function, we ensure that carbon footprint reduction is a central consideration, particularly for applications involving time-sensitive or high-stakes freight such as hazardous materials. This approach aligns with federal and state regulatory frameworks, such as the Clean Air Act, which mandates compliance with stringent greenhouse gas emission standards at the national level (Agency,, 2024), and California’s state-specific policy aiming to reduce emissions by 85% below 1990 levels by 2045 (Institute,, 2024). These regulations underscore the critical need to integrate emissions constraints into optimization models to ensure legal compliance while advancing broader environmental objectives.
Finally, the model’s ability to balance risk aversion and operational efficiency sets it apart. By combining expected cost minimization with CVaR as a risk measure, our approach provides decision-makers with the flexibility to accommodate diverse risk preferences while ensuring robust performance under varying scenarios. This integration of risk and cost efficiency underscores the model’s practical relevance for designing sustainable and resilient intermodal transport systems.

### v2

@@ 1 Introduction
Intermodal freight transportation involves the movement of freight using at least two transportation modes without changing the load unit throughout the process. This approach leverages the cost-effectiveness and high capacity of modes like rail or sea for long-haul segments while relying on trucks for shorter distances. It typically requires consolidating multiple shipments into containers or single units at facilities such as warehouses or manufacturing points. These consolidated units are then transported to intermodal hubs for shipping and eventually transferred to final intermodal nodes for deconsolidation and final delivery. The scale and importance of intermodal freight are evident in the fact that, in 2021, the United States (US) alone transported an average of 53.6 million tons of freight daily, valued at over \$54 billion (of Transportation Statistics,, 2021), emphasizing it’s significant role in the national economy.
Despite its advantages, intermodal freight transport faces several challenges due to the involvement of multiple stakeholders and decision-makers, each with their own priorities and objectives. In the context of our work, a significant challenge is the modal shift from road to rail transport (El Yaagoubi et al.,, 2022). Integrating these modes efficiently requires overcoming additional handling and time requirements compared to direct road transport (Görçün et al.,, 2024). For longer distances, rail’s high fixed costs are offset by increased efficiency, while road transport becomes less cost-effective as distance increases (Zgonc et al.,, 2019).
@@
Another critical challenge is the uncertainty in market demand and spot capacities at intermodal hubs. Demand fluctuations can arise from various factors, including booking changes, cancellations, unforeseen events, and seasonal variations (Wang and Meng,, 2021). Compounding this uncertainty is the stochastic nature of spot capacities, which depend not only on competing cargo from other companies but also on the variability of loading and unloading activities at each station (Sun and Schonfeld,, 2016). These dynamics create a complex environment where maintaining operational efficiency requires careful planning and adaptability.
@@
A key aspect influencing spot capacity allocation is the role of long-term contracts, which are agreements between transport providers and large freight forwarders, ensuring a guaranteed capacity over a time period. Such contracts offer stable and predictable volumes for the transport providers while securing regular service for the customers. However, any unused capacity are often redirected to the spot market. This cascading allocation process introduces uncertainty for spot customers, as the availability of capacity depends on how much of the allocated slots long-term contract holders utilize (Wang and Meng,, 2021). For instance, if contract holders underutilize their slots, the surplus is released for spot demand, but this availability may still be constrained by the need to accommodate delayed or overbooked cargo from other services or other long-term contractors. This interplay between long-term contracts and spot demand underscores the need for flexible and adaptive strategies to ensure optimal utilization of available capacity while addressing the inherent uncertainty in both supply and demand (Lu et al.,, 2020; Delbart et al.,, 2021).
While previous studies have addressed aspects of uncertainty management, facility location, or sustainable logistics separately (Guo et al.,, 2020; Hernandez et al.,, 2024), an integrated framework that jointly considers stochastic demand, spot capacity fluctuations, and carbon emission reduction remains limited. This gap is further discussed in Section 2. Moreover, many existing models assume perfect information or static decisions, overlooking the need for adaptive strategies in real-world intermodal operations.
In this study, we develop a two-stage stochastic optimization framework that explicitly incorporates uncertainty in both demand and train capacities while embedding sustainability considerations into intermodal freight planning. Our model leverages a Conditional Value-at-Risk (CVaR) approach to manage the risks associated with extreme cost outcomes, enabling decision-makers to balance operational costs with risk aversion preferences. In addition, the model penalizes excess carbon emissions, aligning the optimization objectives with broader environmental goals.
To demonstrate the practical relevance of the proposed framework, we apply it to a case study involving major freight rail corridors in the United States. Using real-world data on train schedules, intermodal hub locations, and supply facility distributions, the case study evaluates the model’s performance under multiple uncertainty scenarios and explores the trade-offs between cost, risk, and sustainability objectives.
@@ 2 Literature review
Intermodal freight transport has been widely studied for its potential to improve logistics efficiency, sustainability, and resilience. Recent advancements in optimization techniques, data-driven decision-making, and sustainable transport policies have addressed key challenges such as managing uncertainties, coordinating operations across modes, and minimizing emissions. This section highlights significant recent contributions in these areas, examines emerging approaches, and identifies research gaps, providing context for our proposed model’s contributions.
@@ 2.1 Managing Uncertainties in Intermodal Transport
Managing uncertainties in intermodal transport has been a critical in improving logistics efficiency. Guo et al., (2020) tackled demand fluctuations using a multistage stochastic programming model with a rolling horizon approach, outperforming traditional methods in cost reduction and allocation performance. Likewise, Sun, (2020) applied a fuzzy multi-objective routing model for hazardous materials, using triangular fuzzy numbers and soft time windows to balance risk, cost, and service reliability.
Facility location and disruption management have been used to address transport stochasticity. Badyal et al., (2023) proposed a two-stage stochastic model for terminal location and freight distribution under disruption risks, emphasizing the need of robust planning. Bagheri et al., (2024) compared deterministic, stochastic, and robust models for dry port placement, highlighting the flexibility and environmental benefits of robust optimization under dynamic conditions.
Routing and network design remain pivotal in handling operational variability. Lu et al., (2020) introduced fuzzy programming to address uncertainties in road-rail transport, improving performance through defuzzification and sensitivity analysis. Yang et al., (2024) developed a bi-level genetic algorithm for multimodal routing, effectively minimizing costs and meeting time-sensitive delivery constraints. In addition, Wang et al., (2024) applied adaptive differential evolution to marine transport planning, significantly cutting delays and cost in multimodal networks.
Finally, decision-making frameworks offer a strategic lens for managing systemic uncertainty. Gandhi et al., (2024) used a hybrid multi-criteria framework to assess modal shift risks and enablers, underscoring the role of policy and partnerships. Similarly, Raad and Rajendran, (2024) proposed a fuzzy stochastic model for closed-loop dry port networks, boosting resilience through robust, scenario-based planning. These contributions underscore the growing importance of integrating analytical tools and long-term strategies to build agile, sustainable intermodal systems.
@@ 2.2 Operational Synchronization in Road-Rail Intermodal Transport
Effective synchronization between road and rail is vital for reducing delays and enhancing intermodal efficiency. Newman and Yano, (2000) pioneered train schedule optimization using decomposition techniques to streamline operations and meet delivery timelines. Building on this, Liu et al., (2024) tackled port-side delays with a flexible yard crane scheduling model, optimizing task sequencing and significantly reducing truck congestion for high rail-volume.
Advances in scheduling and routing have improved network performance. Chen et al., (2016) addressed uncertainties in vehicle dispatching at terminals, proving better logistics coordination and fewer delays. Chen et al., (2022) developed a bi-level model integrating route planning and timetable adjustment, showing higher capacity use and lower costs. For emergency scenarios, Tong et al., (2024) introduced a hybrid flow shop model with a memetic algorithm to minimize congestion and transshipment delays, highlighting the value of adaptive scheduling in multimodal systems.
Strategic planning and infrastructure upgrades also play a central role. Delgado et al., (2021) proposed a decentralized terminal management framework, lowering costs and improving planning in Portugal’s rail-road network. Li et al., (2022) used a network equilibrium model to identify bottlenecks on China-Europe routes, guiding infrastructure expansion. In contrast, Kumar and Anbanandam, (2020) highlighted how underdeveloped infrastructure and limited transshipment technologies constrain emerging economies, advocating for targeted investments and policy support.
Finally, addressing systemic inefficiencies is key to sustained coordination. Ke, (2022) proposed robust optimization strategies for hazardous material transport, including rerouting, yard repairs, and third-party carriers, boosting network reliability. Ahmady and Eftekhari Yeghaneh, (2022) added a multi-objective model balancing costs, efficiency, and infrastructure criticality advancing the broader goal of synchronized, resilient intermodal operations.
@@ 2.3 Decarbonization in Intermodal Freight Transport
Decarbonizing intermodal freight transport has driven a wave of innovation through optimization models, technology, and policy. Cheng and Tan, (2019) introduced a multimodal path optimization model that minimized emissions and costs by integrating carbon taxation and trading into time-constrained freight systems. Their genetic algorithm-based solution proved effective in maintaining timeliness while reducing carbon output. Likewise, Sun et al., (2023) developed a fuzzy nonlinear programming model to optimize carbon-efficient road-rail routing, balancing environmental and operational goals under uncertainty.
Technological innovation and alternative fuels are central to emission reduction. Hernandez et al., (2024) assessed biodiesel, e-fuels, hydrogen, and electric locomotives for rail decarbonization, finding electric systems could cut emissions by 46% under favorable costs. Matsuyama et al., (2024) showed that, in Japan’s hinterland transport, emerging technologies could achieve up to 87% emission reductions, surpassing the impact of modal shift policies alone.
Policy-driven modal shifts also show significant promise. In Brazil, Nassar et al., (2023) used a System Dynamics model to stress that infrastructure and policy must advance in tandem for meaningful decarbonization. These studies reinforce the role of integrated policy frameworks in supporting sustainable freight. Finally, consolidation and network optimization contribute meaningfully to decarbonization. Goodarzi et al., (2024) proposed a two-stage stochastic model accounting for disruption scenarios, showing strong environmental and cost benefits. By enhancing network reliability and mitigating disruptions, their work highlighted the value of resilience-focused strategies.
@@ 2.4 Our Contribution
Intermodal freight transport, particularly in road-rail systems, faces significant challenges due to fluctuating demand, constrained spot capacities, and fixed train schedules. To address these complexities, we propose a two-stage stochastic optimization model that supports robust first-stage container preparation decisions, followed by adaptive second-stage recourse actions. This structure offers the flexibility needed to operate efficiently in dynamic, uncertain environments (Tolooie et al.,, 2020; Huang,, 2021).
Our work advances the literature by unifying key priorities, uncertainty management, operational synchronization, and decarbonization into a single integrated framework. Unlike prior studies that examine these elements in isolation (Guo et al.,, 2020; Sun et al.,, 2023), our model also leverages CVaR to account for extreme cost scenarios, striking a deliberate balance between risk resilience, cost efficiency, and environmental impact.
A central innovation of our approach lies in its dynamic adaptation to real-time disruptions, such as variable train capacities and demand shifts. By incorporating spot market capacities at intermodal hubs, our model enables responsive container allocation and preparation within tight operational windows. This marks a departure from static planning models (Badyal et al.,, 2023), allowing decisions to adjust fluidly with on-the-ground realities.
Equally important is our explicit integration of emissions constraints. While previous studies explore low-carbon strategies (Hernandez et al.,, 2024), our model embeds carbon penalties directly into the objective function, ensuring sustainability is not an afterthought but a driving force. This aligns with regulatory imperatives like the Clean Air Act (Agency,, 2024) and California’s aggressive emissions targets (Institute,, 2024), reinforcing the relevance of our work to real-world policy goals.
By combining operational adaptability, environmental measures, and risk-aware planning into a unified stochastic framework, we offer a practical, forward-looking tool for designing resilient and sustainable intermodal systems. The following section outlines the operational context that grounds and motivates our model development.